OpenStudy (anonymous):

HELPPPPPP!!!!!!! Find the value of this number. 5!

OpenStudy (mathstudent55):

This is a factorial number. n! = n * (n - 1) * (n - 2) *...* 2 * 1 Examples: 3! = 3 * 2 * 1 = 6 4! = 4 * 3 * 2 * 1 = 24 Now try 5! =